# It’s Morphing Time: Unleashing the Potential of Multiple LLMs via Multi-objective Optimization

### v1

@@ Abstract
In this paper, we introduce a novel approach for large language model merging via black-box multi-objective optimization algorithms. The goal of model merging is to combine multiple models, each excelling in different tasks, into a single model that outperforms any of the individual source models. However, model merging faces two significant challenges: First, existing methods rely heavily on human intuition and customized strategies. Second, parameter conflicts often arise during merging, and while methods like DARE [1] can alleviate this issue, they tend to stochastically drop parameters, risking the loss of important delta parameters. To address these challenges, we propose the MM-MO method, which automates the search for optimal merging configurations using multi-objective optimization algorithms, eliminating the need for human intuition. During the configuration searching process, we use estimated performance across multiple diverse tasks as optimization objectives in order to alleviate the parameter conflicting between different source models without losing crucial delta parameters. We conducted comparative experiments with other mainstream model merging methods, demonstrating that our method consistently outperforms them. Moreover, our experiments reveal that even task types not explicitly targeted as optimization objectives show performance improvements, indicating that our method enhances the overall potential of the model rather than merely overfitting to specific task types. This approach provides a significant advancement in model merging techniques, offering a robust and plug-and-play solution for integrating diverse models into a unified, high-performing model.
@@ 1 Introduction
Large language models (LLMs) have shown great performance on tasks in various domains such as natural language processing [2], computer vision [3] etc. With the great effort of world-wide contributors in community, a large number of general-purpose pre-trained and task-specific fine-tuned language models have been proposed and made publicly available. However, LLM pre-training or fine-tuning is non-trivial and requires a lot of effort and financial budget. Recently, model merging (MM) has attracted many researchers’ attention. By combining multiple LLMs into a single model with better performance and adaptability on more tasks, MM offers a novel cost-efficient way of obtaining new powerful language models without performing additional training [4, 5], just like the Power Rangers merge their Zords together to form a mighty Megazord [6]. Ideally, MM is supposed to inherit and amplify the strengths from its source models while ignoring their weaknesses. Therefore, the obtained model will be able to tackle the union set of all the tasks where the source models are pre-trained/fine-tuned with better performance. Yet this is achieved without training, which saves a large amount of calculation/financial budget. With the help of open-source toolkits such as mergekit [7, 8], MM has become popular for LLM developing and shown great potential on the Open LLM Leaderboard [9]
However, model merging requires the model maker to have profound knowledge or intuition. automatically discover more capable MM recipes is still in its infancy. To the best of our knowledge, The only related work is [4], where Akiba et al. proposed to use evolutionary algorithms (EAs) to generate powerful model merging recipes which operates in both parameter space and data flow space. However, the approach of using diverse task scores and optimizing the parameter space to explore the comprehensive potential of the final model has been largely overlooked.
@@
In this work, we proposed to leveraging multi-objective optimization for automatic model merging. Specifically, we reformulate the problem of merging multiple models as a multi-objective optimization problem. This automated model merging approach offers several advantages over retraining a new model from scratch: Every model has its unique strengths and weaknesses based on its training dataset and architecture. Merging complementary models into one hyper-efficient system results in a performance powerhouse that leverages the best qualities of each model. The contributions of this work can be summarized as follows:
(1) Automated Model Merging: We introduce a novel approach to model merging that automates the search for optimal merging configurations using multi-objective black-box optimization algorithms, eliminating the reliance on human intuition and customized strategies.
(2) Enhanced Model Potential: Our experiments demonstrate that the proposed method not only preserves individual capabilities while mitigating parameter interference but also improves performance on problem types not explicitly targeted as optimization objectives. This indicates that our method enhances the overall potential of the model, providing a robust and versatile solution for integrating diverse task-specific models into a high-performing, cohesive model.
(3) Comprehensive Evaluation: We validate the effectiveness of our MM-MO method through extensive evaluations on three different types of natural language tasks. The results show that the models merged using our proposed approach achieve superior performance compared to existing methods, highlighting the advantages of our method.
@@ 2.1 Model Merging
One common approach is simply averaging the weights of different models, which has shown promising performance. Model soups [13] achieved notable performance improvements on computer vision tasks. Task arithmetic [14] merges large language models (LLMs) by performing arithmetic operations on task vectors (also known as delta parameters), which represent the differences between the weights of fine-tuned LLMs and the original pre-trained model. Fisher-Weighted Averaging [15] is based on the Laplace approximation, where each model’s posterior is approximated as a Gaussian distribution, with its precision matrix corresponding to its Fisher information.
@@
2.2 Relationship between MM and Multi-Task Learning and Ensemble Learning
Multitask learning (MTL) is a learning mechanism that improves learning efficiency and performance by learning multiple tasks simultaneously. In deep learning, MTL refers to training a neural network to perform multiple tasks by sharing some of the network’s layers and parameters across tasks. MM is similar to MTL since they both try to obtain a machine learning model that can handle multiple tasks when one aim to produce a LLM for different tasks with MM.
Ensemble Learning (EL) Ensemble learning is a machine learning technique that aggregates two or more learners (e.g. regression models, neural networks) in order to produce better predictions. MM is similar to EL since they both try to obtain a machine learning model with multiple source models.
Roughly speaking, MM resembles MTL in final goal (ability on multiple tasks) and EL in input with multiple source models. Other than the above-mentioned points, MM differs from both MTL and EL in that it involves no further training or fine-tuning.
@@ 2.3 Multi-objective Optimization
Multi-objective Optimization (MO) focuses on approximating the PS, targeting at a solution set with good convergence and diversity in the objective space. In this work, we utilize multi-objective optimization to address the complex challenge of model merging. By considering multiple objectives simultaneously, MO enables us to balance various performance metrics, such as different task-specific accuracies. This approach allows us to optimize the trade-offs between different performance criteria, leading to a more robust and effective merged model. Through this methodology, we aim to achieve a unified model that not only excels in specific tasks but also retains a broad and diverse set of capabilities from the individual models. This is particularly important in the context of large language models, where different tasks may require different strengths and features from the base models. By leveraging MO, our method systematically explores the parameter space to identify optimal configurations, ensuring that the final model exhibits both high performance and comprehensive potential across various tasks.
@@ 3.1 Overview
In Figure 1, The dashed lines indicate that the model exhibits relatively superior performance on the corresponding tasks. The final output is the merged model, which integrates the strengths of the individual models and demonstrates superior performance on a wide range of tasks. Our method starts with a collection of source models, each fine-tuned on different tasks. The objective is to merge these diverse models into a single model that not only performs well on the individual tasks but also demonstrates enhanced overall capability across a range of tasks.
The merging process leverages a collection of tasks to guide the optimization. Each source model is evaluated across this task collection to determine its strengths and weaknesses. The task collection includes a variety of tasks, ensuring a comprehensive evaluation of model performance. To improve the ability of the final merged model, we utilize multi-objective black-box optimization. This process involves balancing different performance metrics, such as task-specific accuracy values, while preserving crucial delta parameters. The goal is to achieve a merged model that exhibits both high performance and versatility.
@@ 3.2 Model Merging via Multi-objective Optimization
  • 1)
    Initialization: We start with a collection of pre-trained models, each fine-tuned on different tasks. These models and their configurations serve as the initial candidates for merging.
  • 2)
    Objective Definition: We define multiple objectives that reflect the performance metrics we aim to optimize. In our case, these objectives include task-specific accuracies such as C-EVAL [22] and GSM8K [23] accuracy. These metrics are crucial as they ensure the merged model performs well across different types of tasks.
  • 3)
    Surrogate Model Training: qEHVI leverages surrogate models to approximate the objective functions. We train these surrogate models on the initial set of model configurations and their corresponding performance metrics. This step involves creating a probabilistic model that predicts the performance of unseen configurations.
  • 4)
    Acquisition Function Optimization: The key component of qEHVI is the Expected Hypervolume Improvement (EHVI) acquisition function. This function guides the search by quantifying the expected improvement in the objective space. We optimize this acquisition function to select the most promising model configurations to evaluate next. Specifically, the hypervolume (HV) is a measure of the space dominated by a set of solutions in the objective space, bounded from below by a reference point. The HV indicator is defined as follows:
@@
    $\alpha_{q\text{EHVI}}(\mathcal{X}_{\text{cand}}|\mathcal{P})\approx\hat{\alpha}_{q\text{EHVI}}(\mathcal{X}_{\text{cand}}|\mathcal{P})=\frac{1}{N}\sum_{t=1}^{N}\text{HVI}(\tilde{\boldsymbol{f}}_{t}(\mathcal{X}_{\text{cand}})|\mathcal{P}),$ (3)
    where $\tilde{f}_{t}\sim p(\boldsymbol{f}|\mathcal{D})$ for $t=1,...,N$ and $\mathcal{X}_{\mathrm{cand}}=\{x_{i}\}_{i=1}^{q}$ [21].
  • 5)
    Parallel Evaluation: Unlike traditional methods, qEHVI supports parallel evaluations, which significantly speeds up the optimization process. In each iteration, we select a batch of model configurations and evaluate their performance on the defined objectives (e.g., C-EVAL and GSM8K accuracy).
  • 6)
    Updating the Surrogate Model: After evaluating the selected configurations, we update the surrogate models with the new data. This iterative process continues, refining the surrogate models and improving the acquisition function’s accuracy in identifying high-potential configurations.
  • 7)
    Convergence and Final Selection: The optimization process continues until convergence criteria are met, such as a maximum number of iterations or negligible improvement in the objective space. The final model merging configuration is selected based on the best trade-offs between the multiple performance metrics.
@@
4 Experimental Study
4.1 Setup
Source Models
@@ 4.1 Setup
To develop a model with strong comprehensive abilities across various disciplines, we applied our proposed model merging method to a set of source models, including Qwen1.5-7B-Chat [25], Liberated-Qwen1.5-7B (https://huggingface.co/abacusai/Liberated-Qwen1.5-7B), and firefly-qwen1.5-en-7B (https://huggingface.co/YeungNLP/firefly-qwen1.5-en-7b). All these models were fine-tuned from Qwen1.5-7B [25].
Datasets
We chose C-EVAL [22], a Multi-Level Multi-Discipline Chinese Evaluation Suite, to test the comprehensive abilities of our models across different disciplines. Additionally, we employed a variety of datasets to assess the models’ performance across diverse task scenarios. Specifically, we used GSM8K [23] to evaluate the models’ mathematical reasoning abilities and HumanEval [26] to test their code generation capabilities. Moreover, we used a different dataset for BO search to avoid overfitting the test set. Specifically, we use the valid splits of C-EVAL that were not included in the C-EVAL test set, and we use 1% of the GSM8K train set sampled randomly to ensure diversity and to prevent overfitting.
Evaluation Metrics
C-EVAL: This dataset is used for tasks that involve classification, such as true/false questions and multiple-choice questions. Since these types of tasks require selecting the correct option from a set of possible answers, we use the accuracy metric to evaluate performance. Specifically, we calculate the 5-shot accuracy, which measures the model’s accuracy when given five examples for context before making predictions. This approach helps in assessing the model’s ability to learn from a few examples and generalize to similar tasks. GSM8K: This dataset is designed for mathematical problem-solving tasks, which often require generating accurate and precise solutions from scratch. Since these tasks do not have multiple-choice answers but rather require an exact answer, we evaluate performance using zero-shot accuracy. This means we assess the model’s ability to correctly solve problems without any prior examples or fine-tuning, reflecting its inherent capability to handle complex, unseen mathematical tasks. HumanEval: This dataset is used for evaluating code generation tasks. The nature of these tasks requires generating functional and correct code based on a given problem description. We use the pass@1 metric to evaluate performance, which measures the proportion of problems for which the first generated solution passes all test cases. This metric is crucial as it directly assesses the model’s ability to produce correct and executable code on the first attempt, which is a stringent but realistic measure of practical utility in coding applications.
4.2 Results and Analysis
4.2.1 General Performance

### v2

@@ Abstract
In this paper, we introduce a novel approach for large language model merging via black-box multi-objective optimization algorithms. The goal of model merging is to combine multiple models, each excelling in different tasks, into a single model that outperforms any of the individual source models. However, model merging faces two significant challenges: First, existing methods rely heavily on human intuition and customized strategies to tackle multiple tasks. Second, it’s difficult to search for the great model merging configuration in limited evaluations. To address these challenges, we propose a multi-objective optimization based model merging method named MM-MO. The proposed method can automatically search merging configurations for multiple tasks with multi-objective optimization algorithms. Moreover, to obtain high-quality model merging configurations within a limited number of evaluation iterations, we have made several improvements to multi-objective Bayesian optimization specifically for model merging scenarios. First, we introduced a weak-to-strong method to improve the acquisition strategy. Second, we employed Fisher information to select configurations, further increasing the chances of discovering superior model merging configurations. Third, we designed a sparsity metric as an additional optimization objective to enhance the model’s generalization performance across different tasks. We conducted comprehensive experiments with other mainstream model merging methods, demonstrating that our method consistently outperforms them. Moreover, performance improvements are observed even on the tasks not explicitly targeted as optimization objectives, indicating that our method enhances the overall potential of the model. The proposed method provides a significant advancement in model merging techniques, offering a robust and efficient solution for integrating diverse models into a unified, high-performing model without additional training data.
@@ 1 Introduction
Large language models (LLMs) have shown great performance on tasks in various domains such as natural language processing [1], computer vision [2] etc. With the great effort of world-wide contributors in community, a large number of general-purpose pre-trained and task-specific fine-tuned language models have been proposed and made publicly available. However, LLM pre-training or fine-tuning is non-trivial and requires a lot of effort and financial budget. Recently, model merging (MM) has attracted many researchers’ attention. By combining multiple LLMs into a single model with better performance and adaptability on more tasks, MM offers a novel cost-efficient way of obtaining new powerful language models without performing additional training [3, 4], just like the Power Rangers merge their Zords together to form a mighty Megazord [5]. Ideally, MM is supposed to inherit and amplify the strengths from its source models while ignoring their weaknesses. Therefore, the obtained model will be able to tackle the union set of all the tasks where the source models are pre-trained/fine-tuned with better performance. Yet this is achieved without training, which saves a large amount of calculation/financial budget. With the help of open-source toolkits such as mergekit [6, 7], MM has become popular for LLM developing and shown great potential on the Open LLM Leaderboard [8].
However, model merging requires the model maker to have profound knowledge or intuition. Automatically discovering more capable MM recipes is still in its infancy. To the best of our knowledge, The most related work is [3], where Akiba et al. proposed to use evolutionary algorithms (EAs) to generate powerful model merging recipes which operates in both parameter space and data flow space. However, the approach of using diverse task scores and optimizing the parameter space to explore the comprehensive potential of the final model has been largely overlooked.
@@
In this work, we proposed to leverage multi-objective optimization for automatic model merging. Specifically, we formulate the problem of merging multiple models as a multi-objective optimization problem. This automated model merging approach is able to obtain a performance powerhouse that leverages the best qualities of each model, inheriting their respective strengths of all source models. Furthermore, we validate the effectiveness of our MM-MO method through extensive evaluations across a diverse range of LLM evaluation tasks. The contributions of this work can be summarized as follows:
(1) Automated Model Merging via multi-objective optimization: We introduce a novel approach to model merging that automates the search for optimal merging configurations using multi-objective black-box optimization algorithms, eliminating the reliance on human intuition and customized strategies.
(2) Enhanced Acquisition Strategy: To achieve high-quality model merging configurations with limited evaluation iterations, we introduced a weak-to-strong method to enhance acquisition function that allows previously evaluated superior configurations to guide new configurations’ searching. Besides, we used Fisher information to screen these configurations to increase the chance of discovering superior merging configurations.
(3) Additional Optimization Objective: In addition to task metrics, we have introduced a sparsity metric as an another optimization objective to enhance the model’s generalization performance across different tasks.
The remainder of this paper is organized as follows. Section 2 surveys the existing methods in model merging and multi-objective optimization. Section 3 introduces a novel multi-objective Bayesian optimization algorithm, MM-MO, tailored for model merging. In this section, we also present our improvements to the process of searching for new solutions and the design of new optimization objectives specifically for model merging scenarios. Section 4 presents the performance evaluation of the proposed method against related methods. Finally, we draw conclusions in Section 5.
@@ 2.1 Model Merging
One common approach is simply averaging the weights of different models, which has shown promising performance. Model soups [12] achieved notable performance improvements on computer vision tasks. Task arithmetic [13] merges LLMs by performing arithmetic operations on task vectors (also known as delta parameters), which represent the differences between the weights of fine-tuned LLMs and the original pre-trained model. Fisher-Weighted Averaging [14] is based on the Laplace approximation, where each model’s posterior is approximated as a Gaussian distribution, with its precision matrix corresponding to its Fisher information.
@@ 2.2 Multi-objective Optimization
Multi-objective Optimization (MO) focuses on approximating the PS, targeting at a solution set with good convergence and diversity in the objective space. In this work, we utilize multi-objective optimization to address the complex challenge of model merging. By considering multiple objectives simultaneously, MO enables us to balance various performance metrics, such as different task-specific accuracies. This approach allows us to optimize the trade-offs between different performance criteria, leading to a more robust and effective merged model. Through this methodology, we aim to achieve a unified model that not only excels in specific tasks but also retains a broad and diverse set of capabilities from the individual models. This is particularly important in the context of LLMs, where different tasks may require different strengths and features from the base models. By leveraging MO, our method systematically explores the parameter space to identify optimal configurations, ensuring that the final model exhibits both high performance and comprehensive potential across various tasks.
@@ 3.1 Overview
In Figure 1, we show the overall process of model merging based on multi-objective optimization, which is divided into three parts:
On the left is a schematic diagram of the automatic search process for model merging configurations. Our method starts with a set of source models, each obtained by performing SFT on different datasets using the same pre-trained backbone. These models have their own advantages for different task types. We merge these models to obtain a merged model, evaluate the performance of the merged model on different tasks as optimization objectives , and calculate the sparsity metric of the merged model as an additional optimization objective. For simplicity, we will denote $\boldsymbol{f}(\boldsymbol{x,Tj})$ later as $\boldsymbol{f_{j}}(\boldsymbol{x})$, $j\in\{1,...,M-1\}$. MM-MO generates new model merging configurations $\boldsymbol{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},...\}$ based on these objective values, allowing the creation of new merged models. This process is repeated until a preset evaluation iteration limit is reached, ultimately resulting in an optimal model merging configuration. In the middle is the complete process of using multi-objective Bayesian optimization to search for model merging configurations, detailed in Section 3.2. On the right is the step of generating new configurations. We enhanced the acquisition strategy of multi-objective Bayesian optimization by designing a weak-to-strong method to adjust candidate configurations, and we used Fisher information to select the next generation of model merging configurations, detailed in Section 3.3. We also present the pseudocode of the MM-MO framework in the appendix (Section A).
@@ 3.2 Model Merging via Multi-objective Optimization
1) Initialization: We start with a collection of pre-trained models, each fine-tuned on different tasks. These models and their configurations serve as the initial candidates for merging.
2) Objective Definition: We define multiple objectives that reflect the performance metrics we aim to optimize. Specifically, these objectives include task-specific accuracies, such as those from C-EVAL [22] and GSM8K [23], as well as a sparsity metric $\boldsymbol{g}(\boldsymbol{x})$ used in our experiments. These metrics are crucial as they ensure the merged model performs well across different types of tasks.
3) Surrogate Model Training: qEHVI leverages surrogate models to approximate the objective functions. We train these surrogate models on the initial set of model configurations and their corresponding performance metrics. This step involves creating a probabilistic model that predicts the performance of unseen configurations.
4) Acquisition Function Optimization: The key component of qEHVI is the Expected Hypervolume Improvement (EHVI) acquisition function. This function guides the search by quantifying the expected improvement in the objective space. We optimize this acquisition function to select the most promising model configurations to evaluate next. Specifically, the hypervolume (HV) is a measure of the space dominated by a set of solutions in the objective space, bounded from below by a reference point. The HV indicator is defined as follows:
@@
$\begin{split}\alpha_{q\text{EHVI}}(\mathcal{X}_{\text{cand}}|\mathcal{P})&\approx\hat{\alpha}_{q\text{EHVI}}(\mathcal{X}_{\text{cand}}|\mathcal{P})\\ &=\frac{1}{N}\sum_{t=1}^{N}\text{HVI}(\tilde{\boldsymbol{f}}_{t}(\mathcal{X}_{\text{cand}})|\mathcal{P}),\end{split}$ (3)
where $\tilde{f}_{t}\sim p(\boldsymbol{f}|\mathcal{D})$ for $t=1,...,N$ and $\mathcal{X}_{\mathrm{cand}}=\{x_{i}\}_{i=1}^{q}$ [20]. After obtaining candidate configurations using the acquisition function, we adjusted the candidate configurations using the weak-to-strong method. We then selected the configurations based on their Fisher information to ultimately obtain a new generation of model merging configurations. For details, please refer to Section 3.3.
5) Parallel Evaluation: Unlike traditional methods, qEHVI supports parallel evaluations, which significantly speeds up the optimization process. In each iteration, we select a batch of model configurations and evaluate their performance on the defined objectives (e.g., C-EVAL and GSM8K accuracy).
6) Updating the Surrogate Model: After evaluating the selected configurations, we update the surrogate models with the new data. This iterative process continues, refining the surrogate models and improving the acquisition function’s accuracy in identifying high-potential configurations.
7) Convergence and Final Selection: The optimization process continues until convergence criteria are met, such as a maximum number of iterations or negligible improvement in the objective space. The final model merging configuration is selected based on the best trade-offs between the multiple performance metrics.
@@
3.3 Enhanced Acquisition Strategy
To achieve high-quality model merging configurations within a limited number of evaluation iterations, we have further improved the acquisition function of multi-objective Bayesian optimization for the model merging scenario.
Specifically, we designed a weak-to-strong method, which adjusts the candidate configurations obtained using the acquisition function based on already evaluated superior configurations. In addition, we employ Fisher information to perform environment selection on the adjusted solutions, thereby further enhancing the chance of discovering superior model merging configurations.
The details of these two methods will be introduced in the following two sections.
3.3.1 Weak-to-strong Method
The Weak-to-Strong method has previously been utilized in aligning LLMs. EXPO [25] allows for the direct acquisition of a better-aligned model by extrapolating from the weights of an aligned model and its initial SFT checkpoint, eliminating the need for additional training. Inspired by this methodology, we developed a weak-to-strong method to improve the process of searching for new configurations in our model merging scenario.
The weak-to-strong method operates as follows:
1. Selection of Superior Configurations: From the already evaluated configurations, we select a small number (set to 5 in our experiments) of the top-performing configurations.
2. Differential Evolution: We apply Differential Evolution (DE) [26] to the parameters of these selected configurations. DE is a population-based optimization algorithm that creates new candidate solutions by combining the parameters of existing ones, ensuring diversity and exploration of the search space. Specifically, DE generates new candidate solutions using the following formula:
$\mathbf{v}_{i}^{(k+1)}=\mathbf{x}_{r_{1}}^{(k)}+F\cdot(\mathbf{x}_{r_{2}}^{(k)}-\mathbf{x}_{r_{3}}^{(k)})$ (4)
where $\mathbf{v}_{i}^{(k+1)}$ is the new solution vector for the $i$th individual at generation $k+1$. $\mathbf{x}_{r_{1}}^{(k)}$, $\mathbf{x}_{r_{2}}^{(k)}$, and $\mathbf{x}_{r_{3}}^{(k)}$ are randomly selected distinct individuals from the current population, and $r_{1}\neq r_{2}\neq r_{3}\neq i$. The differential mutation factor $F$ is a scaling factor typically between 0 and 1.
3. Stochastic Perturbation: To introduce an element of randomness and avoid local optima, we apply a stochastic perturbation with a probability of 50%. This perturbation replaces some of the parameters in the candidate configurations with those obtained from the DE process, encouraging exploration beyond the immediate neighborhood of the superior configurations.
This method helps in refining the acquisition function’s output by focusing the search around promising areas in the configuration space, thereby increasing the probability of finding high-quality model merging configurations.
3.3.2 Fisher Information
In the context of our model merging scenario, the purpose is to identify high-quality configurations within a limited number of evaluation iterations. To achieve this, we need to refine our surrogate model efficiently. One way to do this is by carefully selecting which configurations to evaluate based on their potential to improve the surrogate model’s accuracy. Fisher information [27] serves as a pivotal tool in this selection process. By prioritizing configurations with lower Fisher information for evaluation, we can enhance the precision of our surrogate model more efficiently.
Fisher information is a measure of the amount of information that an observable variable provides about an unknown parameter, thereby quantifying the expected information gain from data regarding parameter estimation. Mathematically, for a parameter $\theta$ and a probability density function $p(x;\theta)$ , the Fisher information $I(\theta)$ is expressed as:
$I(\theta)=-\mathbb{E}\left[\frac{\partial^{2}\log p(x;\theta)}{\partial\theta^{2}}\right]=\mathbb{E}\left[\left(\frac{\partial\log p(x;\theta)}{\partial\theta}\right)^{2}\right].$ (5)
In our methodology, the Fisher information for each candidate configuration is calculated using the current surrogate model, which is modeled as a Gaussian process [28], a commonly used surrogate model in Bayesian optimization. The detailed process is as follows:
1. Posterior Predictive Calculation: For each candidate configuration, the Gaussian process surrogate model is employed to compute the posterior predictive mean and variance.
2. Fisher Information Approximation: The variance of the posterior distribution is utilized to approximate the Fisher information for each configuration. Configurations with lower Fisher information indicate regions where the surrogate model’s predictions are less certain.
3. Environment Selection: Configurations with lower Fisher information are prioritized for evaluation. This ensures that configurations contributing more significantly to the model’s uncertainty are evaluated first, thereby enhancing the surrogate model’s overall accuracy.
By incorporating Fisher information into the environment selection process, we aim to improve the efficiency of the evaluation process, enhances the accuracy of the surrogate model more rapidly, and increases the possibility of discovering superior model merging configurations within the constraints of limited evaluation iterations.
3.4 Sparsity optimization objective
In addition to evaluating the performance of each merged model across various tasks, we introduce an additional optimization objective: a sparsity metric $\boldsymbol{g}(\boldsymbol{x})$ This metric is designed to quantify the sparsity of the delta parameters between the merged model and the base model. Specifically, $\boldsymbol{g}(\boldsymbol{x})$ is computed by taking the difference between the Merged Model (MM) and the Base Model to obtain the Delta Parameters, and then calculating the L1-Norm of all Delta Parameters. The sparsity metric $\boldsymbol{g}(\boldsymbol{x})$ can be expressed mathematically as:
$f_{M}(\boldsymbol{x})=\boldsymbol{g}(\boldsymbol{x})=\sum_{i=1}^{n}\|\Delta\theta_{i}\|_{1},$ (6)
where $\Delta\theta_{i}$ represents the delta parameters.
The purpose of incorporating this additional optimization objective is to prevent the final merged model from overfitting to specific evaluation tasks, thereby ensuring the merged model exhibits better overall performance and generalization capabilities.
The rationale behind designing this sparsity metric is supported by the findings of Yu et al. in their research on Dare [21], which demonstrated that for LLMs, removing up to 90% of the delta parameters does not significantly degrade performance. In some cases, the drop rate $p$ can even reach 99%, indicating that the supervised fine-tuning introduces substantial redundancy in the model parameters. This suggests that only minor modifications to the base model parameters are needed to unlock the model’s potential and achieve optimal results.
[…]
4 Experimental Setup
4.1 Source Models
@@
To develop a model with strong comprehensive abilities across various disciplines, we applied our proposed model merging method to a set of source models, including Qwen1.5-7B-Chat [29], Liberated-Qwen1.5-7B [30], and firefly-qwen1.5-en-7B [31]. Besides these 7B models, we also experimented with models of different sizes and pre-trained backbones, including Qwen2-1.5B-Instruct [32], SauerkrautLM-1.5b [33], and Qwen2-1.5B-Ita [34]. Please see Table 5 in Section B.2 for their versions and correspondences with pre-trained backbones.
4.2 Datasets
For our study, we divided the usage of datasets into two distinct phases: configuration search and final model evaluation.
In the configuration search phase, we aimed to optimize the model parameters without overfitting to the test datasets. For this purpose, we used the C-EVAL [22] Validation Set, a Multi-Level Multi-Discipline Chinese Evaluation Suite, to ensure the model’s comprehensive abilities across different disciplines were appropriately calibrated. Additionally, to evaluate and optimize the mathematical reasoning abilities of our model, we sampled 1% of the GSM8K [23] Training Set. Since GSM8K does not have a predefined test set, the validation set is typically used for testing. To avoid data leakage, we selected a small, randomly chosen subset from the training set.
For the final evaluation of our models, we employed a range of datasets tailored to assess various capabilities, including C-EVAL [22], MMLU [35], GSM8K [23], HellaSwag [36], HumanEval [37], MBPP [38], WinoGrande [39] and BIG-Bench Hard (BBH) [40].
Among these, C-EVAL and MMLU are multidisciplinary examinations that evaluate the model’s comprehensive abilities; GSM8K assesses the model’s mathematical reasoning abilities; HellaSwag evaluates the model’s commonsense reasoning abilities; HumanEval and MBPP assess the model’s code generation abilities; WinoGrande evaluates the model’s language understanding abilities; and BBH assesses the model’s comprehensive reasoning abilities. Please see Section B.3 for more descriptions.
By utilizing these datasets, we ensured a comprehensive and balanced evaluation of models across a wide range of abilities and tasks.
4.3 Baselines
We have compared MM-MO with 5 state-of- the-art and classical merging methods, including Linear (Model Soups) [12], Task Arithmetic [13], TIES [17], DARE [21] and Model Breadcrumbs [41]. Since some model merging methods can be combined, there are a total of 7 baselines. Please see Section B.4 for more descriptions.
4.4 Evaluation Metrics
We calculate 5-shot accuracy for C-EVAL, 5-shot weighted accuracy for MMLU, 0-shot and 5-shot accuracy for GSM8K (0-shot for 7B model and 5-shot for smaller 1.5B model), 0-shot accuracy for HellaSwag, pass@1 for HumanEval and MBPP, 0-shot accuracy for WinoGrande, and 3-shot accuracy for BBH.
5 Experimental Results and Analysis
5.1 General Performance